OpenStudy (anonymous):

Consider the function f(x) = x - x^3. determine one or more horizontal shifts that will change its form so that no linear term is present.

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

Send x to x-c, collect terms and solve for c such that the coefficient of the linear term is exterminated.

OpenStudy (anonymous):

\[f(x-c)=(x-c)-(x-c)^3\]\[=(c^3-c)+(1-3c^2)x+3cx^2-x^3\]If you let \[c=\pm \frac{1}{\sqrt{3}}\]and sub. in, your linear (x) term will disappear.

OpenStudy (anonymous):

how did you know that?

OpenStudy (anonymous):

Translating a 1D function to the left or right amounts to shifting the argument by a constant. Just look at y=x^2 versus y=(x-1)^2.

OpenStudy (anonymous):

right, so you moved it so that the roots changed

OpenStudy (anonymous):

The question asks for a horizontal shift that will eliminate the linear term.

OpenStudy (anonymous):

ok , the original function has 3 roots. so shouldnt you move it up vertically ? oh wait... by translating it the roots will average out or something

OpenStudy (anonymous):

ok i see what you did, i got confused because it says shifts* plural , but that was clever

OpenStudy (anonymous):

It's just an application of the definition of 'horizontal shift' along with the condition stipulated in the question that the shift is wanted so that the linear term is extinguished.

OpenStudy (anonymous):

It's a bizarre question.

OpenStudy (anonymous):

you know anything about gabriel horn

OpenStudy (anonymous):

the paradox is that there is no bottom to the horn, so in theory you can keep adding say water to it (ideal water) . but the volume is finite

OpenStudy (anonymous):

ideal water (quantum molecular issues aside)

OpenStudy (anonymous):

It is indeed a weird question, but a similar technique can be used to reduce, for example, a quartic equation to one with no cubic term which makes it easier to solve in some ways. Sounds little random but that's just about application of this thing I've seen before.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Same with suppressing a cubic.

OpenStudy (anonymous):

here is another crazy integral. if you find the surface area of y= ln x from 0 to 1 (so its below x axis), and revolve it about y axis , you get a finite surface area (even though y = ln x goes to negative infinity)

OpenStudy (anonymous):

its called depressed equation , but what you did is straightforward :)

OpenStudy (anonymous):

actually there might be another way, buts kind of long. you can use the symmetric polynomial reduction form . so x^3 - (a+b+c) x^2 + (ab +ac +bc)x - abc , where a,b,c are your roots

OpenStudy (anonymous):

yes, depressed.

OpenStudy (anonymous):

so you want ab + ac + bc = 0

OpenStudy (anonymous):

nevermind, you have it :)

OpenStudy (anonymous):

On the bright side, I like discussions on here with people who can do Maths, and don't spam exactly the same question over and over and do nothing themselves - makes a nice change.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

newton can you look over my solution for work problem, im stuck on the density g thing